Integrate the following[Cosec (x^8)cot (x^8) x^7]dxPlzz

Grade:12

1 Answers

Arun
6 years ago
Dear student
Assume cosec8 x = t
Cosec8 x * cot8 x * 8x7 dx = dt
Hence
Integral become as
dt/8
= t/8  + C
=(1/8) Cosec8 x + C
Regards
Arun (askIITians forum expert)